# TSI-Bench: Benchmarking Time Series Imputation

### v1

@@ C.1 Missing Patterns
To explore the imputation algorithms’ performance on different missing patterns, we set up three different modes: point, subsequence, and block. The missingness construction functions are directly available in the Python library PyGrinder [^9] https://github.com/WenjieDu/PyGrinder. The visualization of them is plotted in figures LABEL:fig:point_missing_pattern, LABEL:fig:subseq_missing_pattern, and LABEL:fig:block_missing_pattern for vivid illustration.
@@ C.2 Evaluation Metrics
$\text{MAE}\left(\hat{y},y,m\right)=\frac{\sum_{d=1}^{D}\sum_{t=1}^{T}|(\hat{y}_{t}^{d}-y_{t}^{d})\cdot m_{t}^{d}|}{\sum_{d=1}^{D}\sum_{t=1}^{T}m_{t}^{d}}$
$\text{MSE}\left(\hat{y},y,m\right)=\frac{\sum_{d=1}^{D}\sum_{t=1}^{T}((\hat{y}_{t}^{d}-y_{t}^{d})\cdot m_{t}^{d})^{2}}{\sum_{d=1}^{D}\sum_{t=1}^{T}m_{t}^{d}}$
$\text{MRE}\left(\hat{y},y,m\right)=\frac{\sum_{d=1}^{D}\sum_{t=1}^{T}|(\hat{y}_{t}^{d}-y_{t}^{d})\cdot m_{t}^{d}|}{\sum_{d=1}^{D}\sum_{t=1}^{T}|y_{t}^{d}\cdot m_{t}^{d}|}$
@@ C.4 Downstream Task Design
PhysioNet2012 and Pedestrian are chosen for the classification task because their every sample has a classification label. In PhysioNet2012, each sample has a label indicating if the patient is deceased in ICU, making it an unbalanced binary classification dataset with 13.9% positive samples. Sample classes in Pedestrian correspond to ten locations of sensor placement.
@@ C.5 Adaptation Paradigm
To tailor the time-series forecasting models for the imputation task, TSI-Bench adopts the imputation framework (including the embedding strategy and the training methodology) from SAITS [20] and proposes such an adaptation paradigm that has been demonstrated to be feasible in this work’s extensive experiments.
@@
Please refer to the SAITS paper [20] for the details about how this imputation framework works.
Figure 12: The training methodology from SAITS [20] for the adapted forecasting models.
@@ C.6 Development Environment
All the experiments were conducted on the King’s College London CREATE[72] HPC platform, equipped with 512G RAM and 128 logical cores CPU (AMD EPYC 9554). One NVIDIA A100 80GB GPU is utilized for acceleration for each job, for a total of 32. The software environment is Ubuntu 12.3.0.
To help reproduce our Python environment easily, we freeze the development environment with Anaconda and save it into files for reference, which is publicly available in our GitHub code repository as well.
@@ Appendix D Full Experimental Results
In the process of constructing TSI-Bench, to analyze the imputation and downstream task performance of various imputation algorithms on different datasets and obtain valuable insights, we conduct extensive experiments. Specifically, we explore the imputation effects of 28 algorithms under 5 missing patterns and evaluate the performance of downstream tasks after imputation. It should be noted that, as PhysioNet2012 and PhysioNet2019 inherently contain a high proportion of missing data, these two datasets are not included in the experiments with 50% or 90% missing rates.
@@
BeijingAir ItalyAir PeMS Pedestrian ETTh1 Electricity PhysioNet2012 PhysioNet2019
@@
BeijingAir ItalyAir PeMS Pedestrian ETTh1 Electricity
@@
BeijingAir ItalyAir PeMS Pedestrian ETTh1 Electricity
@@
BeijingAir ItalyAir PeMS Pedestrian ETTh1 Electricity
@@
BeijingAir ItalyAir PeMS Pedestrian ETTh1 Electricity
@@
Table 5 shows The details of the datasets involved in the experiments under the 10% missing data setting. Table 10 shows the experimental results of the evaluated methods under the 10% missing data setting, including size, the results of the three evaluation metrics previously mentioned, and inference time.
@@
The relative details of the datasets are shown in Table 7, and able 12 shows the imputation performance with 90% point missing. This is a relatively extreme scenario and brings difficulties in estimating the missing values.
@@
The details of the preprocessed datasets in this setting is shown as Table 8, and table 13 shows the imputation results with 50% subsequence missing. It could be observed that the imputation error is generally higher than that with 50% point missing. In this condition, the missing values in the subsequence cannot be easily estimated from their adjacent observed data.
@@ D.6 Visualization of Imputation Performance
Figure 13, 14, and 15 depict the imputation examples by different imputation methods. We pick a multivariate time series from the results of each experiment in Appendix D.2, D.4 and D.5 and show imputation results for a single feature of each time series. We can find that most imputation methods can provide reasonable values for the missing components and thus help improve the data quality.
Figure 13: Visualization of imputation performance on the ItalyAir dataset with 50% point missing. Observed values are represented by blue dots, while missing values are indicated with red crosses.
Figure 14: Visualization of imputation performance on the ItalyAir dataset with 50% subsequence missing. Observed values are represented by blue dots, while missing values are indicated with red crosses.
Figure 15: Visualization of imputation performance on the ItalyAir dataset with 50% block missing. Observed values are represented by blue dots, while missing values are indicated with red crosses.
D.7 Experiments on Downstream Tasks
D.7.1 Classification
@@
Pedestrian (50% subseqence missing rate)
@@
Table 15 and  16 show the classification performance without and with imputation. In general, the classification performance can be improved after an imputation process for time series with missing values. This observation is particularly evident when using XGBoost as the classification method on the PhysioNet2012 dataset with a missing rate of 10%.
D.7.2 Regression
@@
Table 17,  18 and 19 show the regression performance without and with imputation under 3 different missing patterns. It can be found that when XGBoost is used as the regression model, using advanced algorithms to impute missing values can improve the regression performance overall. However, for some simple methods (such as Mean), the regression effect after imputation may not improve or may become worse, which also shows to some extent that it is meaningful and valuable to research on the missing value imputation in time series.
D.7.3 Forecasting
@@
Table 20, 21 and 22 show the forecasting performance without or with imputation under different missing patterns and show the same phenomenon when regression is the downstream task, that is, using advanced algorithms to impute missing values can improve the performance of downstream tasks, while using simple methods like Mean may not have a positive impact on forecasting performance.
D.8 The total number of experiments
@@
HPO experiments. For 24 neural networks, we run 100 HPO trials for each of them on each dataset, 24*100*8 = 19,200
Imputation experiments. For 24 neural network models, we run 5 rounds on 8 datasets with 10% point missing, 6 datasets with 50% point missing, 6 datasets with 90% point missing, 6 datasets with 50% subsequence missing, and 5 datasets with 50% block missing: 24*5*(8+6+6+6+5)=3,720. 4*(8+6+6+6+5) = 124 for 4 naive imputation methods.
Downstream experiments. Experiments on the downstream tasks are performed 5 rounds with 4 algorithms on each dataset imputed by 28 methods. The classification task is on PhysioNet2012 with 10% point missing in the evaluation sets and Pedestrian with 10%, 50%, and 90% point missing, and 50% subsequence missing. The regression and forecasting tasks are on ETT_h1 and PeMS with 10% and 50% point missing, 50% subsequence missing and block missing. 5*4*28*(1+4) + 5*4*28*(4+4) + 5*4*28*(4+4) = 11,760

### v2

@@ C.2 Evaluation Metrics
$\text{MAE}\left(\hat{y},y,m\right)=\frac{\sum_{d=1}^{D}\sum_{t=1}^{T}|(\hat{y}_{t}^{d}-y_{t}^{d})\cdot m_{t}^{d}|}{\sum_{d=1}^{D}\sum_{t=1}^{T}m_{t}^{d}},$
$\text{MSE}\left(\hat{y},y,m\right)=\frac{\sum_{d=1}^{D}\sum_{t=1}^{T}((\hat{y}_{t}^{d}-y_{t}^{d})\cdot m_{t}^{d})^{2}}{\sum_{d=1}^{D}\sum_{t=1}^{T}m_{t}^{d}},$
$\text{MRE}\left(\hat{y},y,m\right)=\frac{\sum_{d=1}^{D}\sum_{t=1}^{T}|(\hat{y}_{t}^{d}-y_{t}^{d})\cdot m_{t}^{d}|}{\sum_{d=1}^{D}\sum_{t=1}^{T}|y_{t}^{d}\cdot m_{t}^{d}|},$
@@ C.4 Downstream Task Design
PhysioNet2012 and Pedestrian are chosen for the classification task because every sample has a classification label. In PhysioNet2012, each sample has a label indicating if the patient is deceased in ICU, making it an unbalanced binary classification dataset with 13.9% positive samples. Sample classes in Pedestrian correspond to ten locations of sensor placement.
@@ C.5 Adaptation Paradigm
To tailor the time-series forecasting models for the imputation task, TSI-Bench adopts the imputation framework (including the embedding strategy and the training methodology) from SAITS (Du et al., 2023) and proposes such an adaptation paradigm that has been demonstrated to be feasible in this work’s extensive experiments.
@@
Please refer to the SAITS paper (Du et al., 2023) for the details about how this imputation framework works.
@@ C.6 Development Environment
All the experiments were conducted on the HPC platform, equipped with 512G RAM and 128 logical cores CPU (AMD EPYC 9554). One NVIDIA A100 80GB GPU is utilized for acceleration for each job, for a total of 32. The software environment is Ubuntu 12.3.0.
To help reproduce our Python environment easily, we freeze the development environment with Anaconda and save it into file for reference, which is available in the supplementary material.
Figure 11: The training methodology from SAITS (Du et al., 2023) for the adapted forecasting models.
@@ Appendix D Full Experimental Results
In the process of constructing the TSI-Bench, to analyse the imputation and downstream task performance of various imputation algorithms on different datasets and obtain valuable insights, we conduct extensive experiments. Specifically, we explore the imputation effects of 28 algorithms under 5 missing patterns and evaluate the performance of downstream tasks after imputation. It should be noted that, as PhysioNet2012 and PhysioNet2019 inherently contain a high proportion of missing data, these two datasets are not included in the experiments with 50% or 90% missing rates.
@@
D.1 10% Point Missing
Table 6 shows The details of the datasets involved in the experiments under the 10% missing setting. Table 11 shows the experimental results of the evaluated methods under the 10% missing setting, including size, the results of the three evaluation metrics previously mentioned, and inference time.
@@
BeijingAir ItalyAir PeMS Pedestrian ETT_h1 Electricity
@@
BeijingAir ItalyAir PeMS Pedestrian ETT_h1 Electricity
@@
BeijingAir ItalyAir PeMS Pedestrian ETT_h1 Electricity
@@
BeijingAir ItalyAir PeMS Pedestrian ETT_h1 Electricity
@@
The relative details of the datasets are shown in Table 8, and able 13 shows the imputation performance with 90% point missing. This is a relatively extreme scenario and brings difficulties in estimating the missing values.
@@
The details of the preprocessed datasets in this setting are shown as Table 9, and table 14 shows the imputation results with 50% subsequence missing. It could be observed that the imputation error is generally higher than that with 50% point missing. In this condition, the missing values in the subsequence cannot be easily estimated from their adjacent observed data.
@@ D.6 Visualization of Imputation Performance
We provide the visualization of the imputation examples by different imputation methods on the ETT_h1 and Electricity datasets. Hereby, we display only the imputation results of SAITS in Figures 12 and 13. Results from other imputation algorithms are available under the folder “Imputation comparison” of supplementary material.
Figure 12: Visualization of imputation performance by SAITS on the ETT_h1 dataset.
Figure 13: Visualization of imputation performance by SAITS on the Electricity dataset.
D.7 Visualization of Time Intervals
Figure 14: Density distribution of time intervals across features in PhysioNet2012 dataset.
We conducted a detailed statistical analysis and visualization of the time intervals in the PhysioNet2012 dataset, as shown in Figures 14. It is crucial to emphasize that the uneven time intervals in medical data represent a significant and common issue. As the figure illustrates, features such as heart rate (HR) and diastolic arterial blood pressure (DiasABP) tend to have shorter time intervals, while features like white blood cell count (WBC) and sodium levels (Na) have much longer intervals. This discrepancy arises from the nature of these measurements—vital signs like heart rate can be continuously monitored by instruments, whereas blood tests are inherently spaced out due to the invasive nature of the procedures.
Furthermore, different types of features, such as continuous variables (e.g., HR, DiasABP) versus categorical variables (e.g., Glasgow Coma Scale [GCS]), exhibit significant variability in their measurement intervals. Existing methods like BRITS and GRU-D attempt to address this by incorporating decay mechanisms, where the influence of an observation decreases as the time interval increases. However, these approaches often overlook the inherent regularity of certain medical measurements. Handling unequal time intervals in healthcare time series is not just about filling in the gaps, but about the need to understand the context in data collection and ensuring that imputation methods respect the temporal dynamics of the data.
D.8 Experiments on Downstream Tasks
D.8.1 Classification
@@
Pedestrian (50% subsequence missing rate)
@@
Table 16 and  17 show the classification performance without and with imputation. In general, the classification performance can be improved after an imputation process for time series with missing values. This observation is particularly evident when using XGBoost as the classification method on the PhysioNet2012 dataset with a missing rate of 10%.
D.8.2 Regression
@@
Table 18,  19 and 20 show the regression performance without and with imputation under 3 different missing patterns. It can be found that when XGBoost is used as the regression model, using advanced algorithms to impute missing values can improve the regression performance overall. However, for some simple methods (such as Mean), the regression effect after imputation may not improve or may become worse, which also shows to some extent that it is meaningful and valuable to research the missing value imputation in time series.
D.8.3 Forecasting
@@
Table 21, 22 and 23 show the forecasting performance without or with imputation under different missing patterns and show the same phenomenon when regression is the downstream task, that is, using advanced algorithms to impute missing values can improve the performance of downstream tasks, while using simple methods like Mean may not have a positive impact on forecasting performance.
D.9 The Total Number of Experiments
@@
HPO experiments. For 24 neural networks, we run 100 HPO trials for each of them on each dataset, with $24*100*8=19,200$ experiments.
Imputation experiments. For 24 neural network models, we run 5 rounds on 8 datasets with 10% point missing, 6 datasets with 50% point missing, 6 datasets with 90% point missing, 6 datasets with 50% subsequence missing, and 5 datasets with 50% block missing, with $24*5*(8+6+6+6+5)=3,720$ experiments. We also conduct $4*(8+6+6+6+5)=124$ experiments for 4 naive imputation methods.
Downstream experiments. Experiments on the downstream tasks are performed in 5 rounds with 4 algorithms on each dataset imputed by 28 methods. The classification task is on PhysioNet2012 with 10% point missing in the evaluation sets and Pedestrian with 10%, 50%, and 90% point missing, and 50% subsequence missing. The regression and forecasting tasks are on ETT_h1 and PeMS with 10% and 50% points missing, 50% subsequence missing and block missing. So the total number of downstream tasks is $5*4*28*(1+4)+5*4*28*(4+4)+5*4*28*(4+4)=11,760$ experiments.